# Governance-Aware Vector Subscriptions for Multi-Agent Knowledge Ecosystems

### 21 Mar 2026 v1

@@ 3.3 Properties
Policy soundness. For each dimension $i$ of the policy predicate, if $\textsc{Policy}_{i}$ evaluates to false, the notification is suppressed. This holds independently per dimension: a chunk with $\delta(c)=1$ is never delivered to a marketing agent, regardless of other dimensions. Formally: for all dispatched notifications and for all dimensions $i$, $\textsc{Policy}_{i}(a,c)=\text{true}$.
@@
Curation guarantee. The $\textsc{Curation}(c)$ condition ensures only validated content triggers content subscriptions. This is enforced structurally: content subscription matching is triggered only by the status transition to active. The system also supports review subscriptions (triggered on chunk creation as proposed) for agents that monitor incoming content for review. Each subscription declares a trigger_status (active, proposed, or both), separating knowledge consumption from governance participation.
@@ 5.4 Threats to Validity
Declarative trust. Agents may misdeclare their profiles. This is an inherent limitation of declarative policy systems. We identify three unimplemented mitigation strategies: (1) runtime behavior monitoring, (2) reputation-gated subscription privileges, and (3) periodic automated auditing. The subscription mechanism provides policy soundness given correct inputs; ensuring input correctness is an orthogonal problem requiring complementary mechanisms. This is the primary residual security risk.
@@ 6.2 Relation to Broader Architecture
Governance-aware vector subscriptions are one component of the Cognitosphere, a governance-first architecture for multi-agent knowledge ecosystems. The broader architecture includes deliberative curation (proposal-review-validation lifecycle), reputation systems (dual-track contribution and policing scores), and inter-component governance flows. The subscription mechanism benefits from the curation guarantee but can operate independently.

### 27 Mar 2026 v2

@@ 2.4 Data Handling Policies for Agents
2.5 Normative Multi-Agent Systems
The governance dimension of our work connects to the normative MAS tradition. Electronic institutions esteva2002 formalize agent interactions through structured scenes with permitted speech acts and norm enforcement, providing a theoretical foundation for constrained multi-agent interaction. Boella and van der Torre boella2004 distinguish constitutive norms (rules that create institutional facts, such as “a chunk that passes review counts as validated knowledge”) from regulative norms (rules that constrain behavior, such as policy predicates restricting notification dispatch). Our mechanism embeds both types: the curation condition is constitutive (the status transition to active creates an institutional fact—validated content), while the policy predicate is regulative (it constrains which agents may receive which notifications). Ostrom ostrom1990 identified design principles for commons governance, later formalized for computational agents by Pitt et al. pitt2012 , establishing that self-organizing institutions can sustainably govern shared resources. A curated knowledge base is such a shared resource. Köster et al. koster2022 showed empirically that even arbitrary normative structures improve cooperation among artificial agents, suggesting that governance overhead in agent ecosystems has operational value beyond its direct enforcement function. Recent work has begun bridging the normative MAS tradition to LLM-based agents: Savarimuthu et al. savarimuthu2025 survey how normative reasoning capabilities transfer to large language models, and Kampik et al. kampik2022 map normative MAS governance concepts to web-based autonomous agents.
These frameworks address how agents reason about and comply with norms in transactional interactions. They do not address the specific problem we tackle: composing semantic similarity matching with multi-dimensional policy filtering in a push notification pipeline over curated content. Our work is complementary—it operationalizes normative constraints within a concrete subscription mechanism rather than modeling normative reasoning at the agent level.
@@ 3.3 Properties
Policy soundness. For each dimension $i$ of the policy predicate, if $\textsc{Policy}_{i}$ evaluates to false, the notification is suppressed. This holds independently per dimension: a chunk with $\delta(c)=1$ is never delivered to a marketing agent, regardless of other dimensions. Formally: for all dispatched notifications and for all dimensions $i$, $\textsc{Policy}_{i}(a,c)=\text{true}$. In normative MAS terms, this property corresponds to norm compliance: every dispatched notification satisfies all regulative constraints boella2004 .
@@
Curation guarantee. The $\textsc{Curation}(c)$ condition ensures only validated content triggers content subscriptions. This is enforced structurally: content subscription matching is triggered only by the status transition to active. In normative MAS terms, this transition is a constitutive rule boella2004 : it creates an institutional fact (“this content is validated”) that enables downstream actions (notification dispatch). The system also supports review subscriptions (triggered on chunk creation as proposed) for agents that monitor incoming content for review. Each subscription declares a trigger_status (active, proposed, or both), separating knowledge consumption from governance participation.
@@ 5.4 Threats to Validity
Declarative trust. Agents may misdeclare their profiles—a form of rational norm violation boella2004 where an agent calculates that the benefit of accessing restricted content outweighs the risk of detection. This is an inherent limitation of declarative policy systems. We identify three unimplemented mitigation strategies: (1) runtime behavior monitoring, (2) reputation-gated subscription privileges, and (3) periodic automated auditing. The subscription mechanism provides policy soundness given correct inputs; ensuring input correctness is an orthogonal problem requiring complementary mechanisms. This is the primary residual security risk.
@@ 6.2 Relation to Broader Architecture
Governance-aware vector subscriptions are one component of the Cognitosphere, a governance-first architecture for multi-agent knowledge ecosystems. The broader architecture includes deliberative curation (proposal-review-validation lifecycle), reputation systems (dual-track contribution and policing scores), and inter-component governance flows. The subscription mechanism benefits from the curation guarantee but can operate independently. A companion paper johnson2026paper2 analyzes the governance challenges of agent knowledge curation in depth, mapping platform governance lessons to the agent setting and proposing design considerations for the curation layer that precedes subscription dispatch.